OpenStudy (anonymous):

WILL MEDAL AND FAN The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below, where A = the pounds of Arabica and R = pounds of Robusta: 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds Numerical Answers Expected!

OpenStudy (anonymous):

you got medal me

OpenStudy (anonymous):

HUH

OpenStudy (anonymous):

Hey check on what I sent you :) @A6020072944

OpenStudy (anonymous):

WHAT WAS IT

OpenStudy (anonymous):

CAN ANYBODY EXPLAIN THE PROBLEM TO ME WITH THE ANSWER PLEASE

OpenStudy (jjwalton04):

the equation given is 700A+1200R=1,000,000 why the cost should be fixed at 1,000,000 is beyond me, but anyway that is what it says then it says A=1000 so now the equation is 700×1000+1200R=1,000,000 or 700,000+1200R=1,000,000 which i would solve in two steps 1) subtract 7000,000 to get 1200R=300,000 2) then divide by 1200 to get R=250